OpenStudy (anonymous):

Let 2 = 9^M and 7 = 9^W. What is log subscript 9 of 28? and log subscript 9 of 3.5^1/6 ?

OpenStudy (paxpolaris):

so, \(\large M=\log_9(2)\) and \(\large W=\log_9(7)\)

OpenStudy (anonymous):

Could you explain?

OpenStudy (anonymous):

What is log∨9 of 28?

OpenStudy (paxpolaris):

you get how i got what M and W equal...?

OpenStudy (anonymous):

no not really, could you explain?

OpenStudy (paxpolaris):

it's the definition of logarithm, isn't it...? if \(5^2=25\) .... then \(\large \log_5(25)=2\)

OpenStudy (anonymous):

so did you add the logs to the given information-- how did you get 9 to the power of M as log of 9?

OpenStudy (anonymous):

log ⋄ 9∧M = log 2 ?

OpenStudy (paxpolaris):

it is just the definition of logs if, \(\Large a^m=n\) then, \[\Large \log_a (n) =m\] ... but i think we can show it your way too. start with the original: \(\large 2=9^M\) then, take log base 9 of both sides:\[\large \log_9(2) = \log_9\left(9^M\right)\]

OpenStudy (anonymous):

okay, this is beginning to make more sense--i definitely had trouble with that rule.

OpenStudy (paxpolaris):

\[\large \implies \log_9(2)=M \cdot \log_9(9)\]\[\large \implies \log_9(2)=M \cdot 1\]\[\large \implies \log_9(2)=M \]

OpenStudy (anonymous):

okay that makes sense. thanks.

OpenStudy (paxpolaris):

now you have to write the logs you have in terms of logs of 2 and 7 only...

OpenStudy (paxpolaris):

\[\large {\log_9(28)= \log_9(7\cdot4) \\= \log_9(7)+\log_9(4) \\= \log_9(7)+\log_9(2^2)\\= \log_9(7)+2\log_9(2)\\=W+2M }\]

OpenStudy (anonymous):

wow that is really clear. great explanation!

OpenStudy (paxpolaris):

you can do the 2nd one yourself ... what do you get

OpenStudy (anonymous):

I'm redoing the problem you just showed me for comprehension-- in a few minutes I'll probably be ready for the second equation.

OpenStudy (anonymous):

would putting the equation within a radical help?

OpenStudy (anonymous):

1/6 log 9 (3.5) ?

OpenStudy (paxpolaris):

yes ... keep going

OpenStudy (paxpolaris):

\[\large {\log_9\left( 3.5^{1/6} \right)\\=\frac 16 \log_9(3.5)\\=??}\]

OpenStudy (anonymous):

yes, that's where I am.

OpenStudy (paxpolaris):

3.5 = 7/2 ....

OpenStudy (anonymous):

Ha you're right. It was right in front of me.

OpenStudy (anonymous):

1/6 log (W/M) ?

OpenStudy (anonymous):

1/6 + log W/ log M ?

OpenStudy (paxpolaris):

\[\large {\log_9\left( 3.5^{1/6} \right)\\=\frac 16 \log_9(3.5)\\=\frac 16\left[ \log _9\left( \frac72 \right) \right]}\] you can only replace with W and M when log 7 and log 2 are separate. forget about the 1/6 for now...>>> what's the rule for logs of a division???

OpenStudy (anonymous):

they subtract.

OpenStudy (paxpolaris):

right ... use that to separate log 7 from log 2 \[\log \left( \frac ab \right) = \log (a) -\log (b)\]

OpenStudy (anonymous):

log 9 (7) - log 9 (2) .

OpenStudy (paxpolaris):

\[\large {\frac 16\left[ \log _9\left( \frac72 \right) \right]\\ }\] \[\Large ={\frac 16\cdot\left[ \log _9(7) -\log _9(2)\right]\\ }\]now you can replace with W and M

OpenStudy (anonymous):

1/6 (W-M) . Is that the answer?

OpenStudy (paxpolaris):

\[\Huge \checkmark\]

OpenStudy (anonymous):

Thank you so much.

OpenStudy (paxpolaris):

\[\large {W-M \over6}\]

OpenStudy (anonymous):

Ah. right.

OpenStudy (paxpolaris):

no problem ... it's just fine as you wrote it too

OpenStudy (anonymous):

See you later.